### v1

Abstract
We introduce GraviBERT, a novel deep learning framework for inference of gravitational wave time series, which relies on an Inception-inspired multi-scale convolutional feature extractor with a transformer encoder along a suitable regression head. GraviBERT is trained in two stages: a BERT-style pretraining phase, where the model learns to predict masked segments in feature space in order to capture universal patterns and physics, followed by supervised fine tuning for accurate parameter estimation. This approach demonstrates impressive improvement across multiple metrics compared to training from scratch: On in-domain data, it reduces the mean absolute error for point-estimate parameter inference by up to 30%, and training convergence accelerates by up to $6\times$. Moreover, the mean relative precision of the inferred masses and distances at low SNR reaches the few-percent level, while the mean absolute error in the effective spin about $10^{-3}$. For domain adaptation to new detector noise profiles, the pretrained model demonstrates remarkable efficiency, converging up to $15\times$ faster on small target datasets and reduces estimation errors by up to $\sim 45$%, indicating that it learns sufficient detector-agnostic representations. Cross-approximant transfer demonstrates comparable transferability, achieving up to $44\%$ MAE reductions across all parameters and up to $15\times$ training speedups, with $R^{2}$ scores consistently exceeding $0.9$ for mass parameters at SNR = 10 compared to $0.74$ - $0.87$ when training from scratch. Notably, GraviBERT works directly with the noisy waveforms. The final regression head of the model can be adapted for various downstream tasks after pretraining, positioning it as a step towards foundation-style models in gravitational wave and multi-messenger astronomy.
1 Introduction
Time series analysis lies at the heart of numerous scientific and applied domains, from fundamental physics to finance and climate modeling. Traditionally, architectures such as autoregressive models, convolutional neural networks (CNNs), and recurrent neural networks (RNNs) have been widely employed for this task, with RNNs and their variants such as Long Short-Term Memory (LSTM) networks and Gated Recurrent Units (GRUs) the dominant choice due to their ability to model temporal dependencies in sequential data [Hochreiter1997, Cho2014].
More recently, transformer architectures, originally developed for natural language processing, have redefined the landscape of time series modeling by offering significantly enhanced scalability and parallelism, as well as the ability to capture subtle global dependencies through the so–called self-attention mechanism [vaswani2017attention]. Yet, despite their growing success, the application of transformers to Bayesian parameter inference, particularly in the computationally demanding domain of gravitational wave astronomy, remains largely unexplored. In this work, we introduce a transformer-based architecture tailored for analysis of gravitational waveforms. Our approach bridges a gap in the field, demonstrating that transformers can excel in both efficiency and accuracy in this context, while it represents a step towards the development of foundation models for gravitational wave science which could be used for multi-modal tasks.
The discovery of gravitational waves (GWs) marked a transformative moment in physics, opening an unprecedented observational window into the Universe [Abbott2016]. GWs are ripples of spacetime produced by the acceleration of compact astrophysical objects such as black holes and neutron stars, and are physically recorded by the detector as time-domain, oscillatory signals characterized modulated by the underlying dynamics of the source system. Beyond confirming General Relativity (GR) in the previously untested strong-field regime, GWs enable inference of key physical properties of their sources, such as component masses, spins, orbital dynamics, and geometrical configurations [TheLIGOScientificCollaboration2021]. This makes them invaluable for probing the nature of gravity over astrophysical and cosmological distances, the physics of stellar populations, testing theories related to dark matter or environmental effects, the physics of dark energy, or exotic fields [Berti2015, Saltas:2018fwy, Abac:2025saz, LISA:2022kgy, Barausse:2020rsu]. Reconstructing this wealth of information from noisy observational data constitutes a formidable inverse problem. Aside the challenges arising from the non-linear nature of Einstein’s equations, the low signal-to-noise ratios (SNR) and various instrumental and modelling systematics present hurdles with the waveform reconstruction. Recent work has shown that the computational challenges for higher-order modes may be somewhat mitigated through efficient search strategies [Estelles2024higher, Heinzel2024higher], though the problem remains significant for precessing systems with eccentricity. Consequently, the development of fast, accurate, and robust methods for GW signal analysis is essential to fully exploit the scientific potential of current and next-generation detectors.
Machine learning (ML) has emerged as a promising tool to tackle the above challenge. GW data analysis proceeds in two stages: signal detection via matched filtering [abbott2016observation, allen2005chi2], which identifies candidate events by cross-correlating data with waveform templates, followed by parameter estimation using Bayesian inference [skilling2006nested, veitch2015parameter, ashton2019bilby] for confirmed detections. ML-based approaches can substantially accelerate GW data analysis [^1] We should note here that it is impossible to provide here an exhaustive list of previous works, and for reviews we refer to [cuoco2021enhancing, stergioulas2024machine, torres2024applications, ormiston2024machine, alyasiri2023ai, george2018deep, george2018deep2_detection, aasi2024contribution, huerta2022inference, krastev2020real, zhao2023space].. GW parameter estimation conventionally relies on matched filtering [abbott2016observation, wainstein1962extraction, allen2005chi2] and Bayesian inference [skilling2006nested, veitch2010bayesian, vandersluys2008parameter, cornish2015bayeswave, veitch2015parameter, ashton2019bilby, speri2024probing, cole2025sequential] (see also [thrane2019introduction, meyer2022computational] for reviews), requiring up to weeks per event. Recent neural network approaches achieve dramatic speedups: normalising flows accelerate posterior sampling [williams2021nested, williams2023importance, green2020gravitational, hu2025long], neural posterior estimation enables real-time inference through frameworks like DINGO, which solves the full 15-parameter BBH problem [dax2023dingo, dax2023nature], and related NPE methods [dax2021real, dax2023neural], conditional variational autoencoders reduce inference time from weeks to seconds [gabbard2022bayesian], and neural network emulators enable rapid likelihood evaluation [chua2019reducing, williams2021gaussian]. Recent transformer-based approaches for GWs include work on overlapping signal parameter estimation in Einstein Telescope (ET) scenarios [papalini2025transformer] and pretrained audio transformers [chatterjee2024transformer] (see also [Wang:2025aqk, Chan:2025pdf, Inguglia:2025cig] for similar applications with alternative architectures). Our work differs in its focus on unsupervised masked pretraining directly on noisy GW data for learning generalizable representations and demonstrating transfer learning across detector configurations, rather than addressing a specific astrophysical application. Score-based models [langendorff2023score] and density estimation methods using masked auto-regressive flows [green2021complete] offer additional approaches for robust Bayesian inference. Matched filtering itself presents substantial computational challenges, with template banks that scale prohibitively for higher-order modes and eccentricity in third-generation detectors [owen1999matched], making efficient alternatives critically urgent. However, ML methods typically depend on assumptions about noise, waveforms, and detector modeling, requiring retraining when configurations change. Most methods lack generalizability across signal types or parameter ranges, and their black-box nature can obscure uncertainty quantification and systematic errors.
ML methods come with limitations – they usually depend on assumptions about the underlying noise, waveform or detector modelling. Training must be typically repeated whenever detector configurations, noise characteristics, or theoretical models evolve or change. In this regard, most methods lack a fair amount of generalizability across different signal types or parameter ranges, creating brittleness in operational deployment. Often, the black-box nature of these approaches obscures uncertainty quantification and makes systematic error detection challenging.
Transformers represent a revolutionary architecture that fundamentally transformed natural language processing [vaswani2017attention]. Their distinctive self-attention mechanism enables dynamic weighting of different temporal components within sequences while processing data in parallel, facilitating both computational scalability and the capture of long-range dependencies that traditional recurrent architectures struggle to model effectively. The architecture produces rich encoded representations (embeddings) that serve as versatile foundations for diverse downstream tasks through appropriate task-specific heads. This capability to generate transferable representations lies at the core of foundation models and enables multi-modal applications across domains. Pioneering examples of such transformer-based foundation models include BERT [Devlin2018] for bidirectional language understanding and the GPT series [Radford2018, Radford2019] for autoregressive generation, both of which have demonstrated remarkable versatility and performance across numerous tasks through pretraining and fine-tuning paradigms.
While full Bayesian posterior estimation remains the gold standard for final scientific analysis [veitch2015parameter, ashton2019bilby], our work focuses on demonstrating for the first time the effectiveness of transformer-based architectures with unsupervised pretraining for GW parameter inference. We use point estimates with MC Dropout uncertainty quantification [gal2016dropoutbayesianapproximationrepresenting] as a computationally efficient framework to evaluate our architectural and training novelties. Extension to full Bayesian inference, for example, by replacing the regression head with normalising flows—represents a natural but separate direction for future work. Our primary contributions therefore are: (i) demonstrating that transformer-based architectures effectively capture GW features, (ii) showing that unsupervised pretraining significantly improves training efficiency and accuracy, and (iii) establishing that pretrained models present with promising transfer learning capabilities. The name "GraviBERT" reflects our model’s conceptual inspiration from BERT’s masked pretraining approach, though adapted for continuous-valued time series rather than discrete tokens.
We structure the paper as follows: In Section 2 we discuss our choice of waveform modelling and data generation. In Section 3 we introduce our model architecture, and proceed discussing its training in Section 4. In Section 5 we present our results and discuss our conclusions in Section 6.
2 Modelling and data generation
The GW signal from a binary merger, such as that of two black holes or neutron stars, exhibits three distinct evolutionary phases: inspiral, merger, and ringdown [Maggiore2008, Blanchet2014]. During the inspiral phase, the two compact objects orbit each other at relatively large separations, gradually losing orbital energy through gravitational radiation emission. As the objects approach closer separations, their gravitational interaction becomes highly nonlinear, marking the transition to the merger phase. Finally, the merger culminates when the two bodies coalesce into a single object, producing the characteristic peak amplitude in the observed waveform, followed by the ringdown phase where the remnant object settles to equilibrium [Pretorius2005, Campanelli2006]. The physical characteristics of these waveforms are essential for our subsequent analysis. At leading order in the post-Newtonian expansion, the GW strain can be approximated as [Blanchet2014]:
$h(t)\simeq\frac{\mathcal{M}_{c}^{5/3}}{d}\,\omega(t)^{2/3}\cos\left[2\Phi(t)\right],$ (1)
where $\mathcal{M}_{c}\equiv(m_{1}m_{2})^{3/5}/(m_{1}+m_{2})^{1/5}$ is the chirp mass, $\omega(t)$ represents the center-of-mass orbital frequency, $\Phi(t)$ denotes the orbital phase, and $d$ is the luminosity distance to the source. The orbital phase evolution is governed by [Blanchet2014, Buonanno2009]:
$\frac{d\Phi}{dt}\propto\left(\frac{G\mathcal{M}_{c}}{c^{3}}\right)^{-5/8}\cdot(t_{c}-t)^{-3/8}\cdot\left[1+\cdots+\chi v^{3}+\mathcal{O}(v^{n})\right],$ (2)
where $t_{c}$ denotes the coalescence time, $v\equiv\left(G(m_{1}+m_{2})\omega\right)^{1/3}$ is the orbital velocity, and $\chi_{\rm{eff}}\equiv(m_{1}s_{1}+m_{2}s_{2})/(m_{1}+m_{2})$ represents the effective spin parameter.
This formulation reveals several critical features for parameter estimation: (i) The luminosity distance provides an overall amplitude suppression scaling as $\sim 1/d$; (ii) At leading post-Newtonian order, the phase evolution is primarily governed by the binary’s masses; (iii) Higher-order corrections introduce mass-spin coupling terms that create parameter degeneracies during inference [Cutler1994, Vallisneri2008]. These degeneracies, particularly between mass and spin parameters, pose significant challenges for neural network-based parameter estimation and must be carefully addressed in our architecture design.
We utilise the Effective One Body (EOB) formalism through the SEOBNRv4 approximant [Bohe2017], which delivers accurate modeling of both inspiral and merger phases through a combination of analytical post-Newtonian theory and numerical relativity calibration. We will focus on binary black hole (BBH) mergers with mass ratios $q=m_{2}/m_{1}$ ranging between $q\sim 0.1-1$. These systems fall within the expected validity of our approximant and lie within the detectable range of current and next-generation GW observatories, such as Advanced LIGO and the ET.
GWs manifest as time-series strain measurements in interferometric detectors. Given the limited number of confirmed detections from current observatories [Abbott2021], training robust neural networks necessitates extensive simulated datasets. For waveform generation, we employ the open-source PyCBC library [Nitz2024], which provides access to state-of-the-art waveform approximants. Our training dataset is constructed using a grid-based approach over source masses, spins and luminosity distance. We exclude source orientation or other parameters from our grid, as we are here interested in a first proof-of-concept analysis. We employ a fixed sampling rate of 4096 Hz with a lower frequency cutoff of $f_{\rm low}=20$ Hz and simulate ET detector noise using as our reference the ET-B power spectral density (PSD) [Hild2011]. While ET typically uses $f_{\rm low}=5$ Hz to maximize sensitivity, we use 20 Hz for this proof-of-concept to reduce computational costs and waveform duration, noting that the lower cutoff primarily affects low-mass systems and does not fundamentally limit the applicability of our method.
The duration of each waveform depends on the source parameters, with more massive systems producing shorter signals. We write,
$h=h\left[p_{i};t\right],\quad p_{i}=\{m_{1},m_{2},\chi_{\text{eff}},d\},$ (3)
with the effective spin $\chi_{\rm eff}$ defined as $\chi_{\rm eff}\equiv(m_{1}s_{1}+m_{2}s_{2})/(m_{1}+m_{2})$, where $m_{1}$ and $m_{2}$ are the component masses (with $m_{1}\geq m_{2}$ by convention), and $s_{1}$ and $s_{2}$ are the dimensionless spin parameters aligned with the orbital angular momentum, and the parameter ranges
$m_{1},m_{2}\in[5,150]~M_{\odot},\quad s_{1},s_{2}\in[-0.4,0.4].$ (4)
The distance $d$ denotes an effective luminosity distance used to rescale the signal amplitude and enforce a fixed target SNR. When detector noise is included, $d$ is adjusted for each waveform realization using the scaling $h(t)\propto 1/d$. We focus on two fixed SNR values (10 and 30) to enable controlled comparison of model performance under different (conservative) noise conditions. This assumption must be compared with the continuous SNR distributions expected in realistic astrophysical observations. Extension to variable SNR distributions motivated by astrophysical populations is a direction for future work.
We model single-channel strain data compatible with either the L-shaped configuration or a single arm of the triangular configuration. The full triangular ET design provides three independent interferometric data streams, which would require multi-channel architectures for optimal analysis. For this proof-of-concept, we focus on single-channel analysis to isolate the benefits of our pretraining strategy and an extension to multi-channel processing for the complete triangular configuration is straightforward and remains for future work. We generate the noise realization $n(t)$ by colouring Gaussian noise based on the given PSD and construct the observed signal (s) as
$s(t)=h(t)+n(t).$ (5)
A sample of our waveforms in frequency space is shown in Figure 1.
Figure 1: The ET’s power spectral density (black, ET-B) overlaid with the frequency-domain representations of selected waveforms from our sample at fixed SNR = 30. Different waveforms intersect the PSD at different stages of their evolution. Our analysis uses waveforms with a low-frequency cutoff at $f_{\rm low}$ = 20 Hz, as explained in Section 2.
We emphasise that we focus on the four intrinsic parameters and do not model extrinsic parameters such as sky location ($\alpha,\delta$), inclination angle ($\iota$), polarization angle ($\psi$), or coalescence phase ($\phi_{c}$). Consequently, we do not apply the detector antenna pattern function $F^{+}(\alpha,\delta,\psi,t)$ and $F^{\times}(\alpha,\delta,\psi,t)$ that would project the GW polarizations onto the detector frame. Our strain data represents the waveform $h(t)$ in the source frame at optimal orientation, rather than the detector-frame strain $s(t)=F^{+}h_{+}(t)+F^{\times}h_{\times}(t)$ that would be observed in a real interferometer. This simplification significantly reduces the complexity of the parameter space but limits direct applicability to observational data. However, it serves our primary objective: demonstrating that unsupervised pretraining and transfer learning are effective strategies for GW analysis. The objective of this work are orthogonal to whether the full parameter space is modelled, that is, the same pre-training approach can be applied to architectures handling all 15 parameters.
We note that the $(m_{1},m_{2})$ parameterization used here exhibits stronger correlations than alternative parameterizations such as chirp mass and mass ratio. Future implementations could benefit from using $({\cal M}_{c},q)$ or similar parameterizations that align more naturally with the physical variations encoded in the waveform, potentially improving training efficiency and reducing parameter degeneracies.
We generate training data on a regular grid over the parameter space rather than drawing randomly from astrophysically motivated priors. While production systems typically use prior-based sampling, grid-based generation offers advantages for our proof-of-concept study: (i) it ensures uniform coverage of the parameter space, (ii) it enables systematic evaluation of model performance across parameter ranges, and (iii) it provides a controlled setting for isolating the effects of our pretraining and transfer learning strategies. We emphasise that the network performs regression on continuous values, i.e the grid structure in training data does not reduce this to a classification problem, as the network must interpolate between grid points and is trained with mean squared error loss on continuous targets. This represents a simplification compared to realistic inference scenarios.
3 Model architecture
3.1 Overview
Our model architecture, schematically shown in Figure 4, is structured as a sequential combination of three primary components: i) a convolutional neural network (CNN) feature extractor, ii) a transformer encoder, and iii) a regression multi-layer perceptron (MLP) head. This hybrid architecture is the core strength of our model. The model combines a CNN feature extractor ($\mathcal{F}$) with a Transformer encoder ($\mathcal{T}$), leveraging the complementary strengths of both architectures for analysing time-series data [foumani2023deeplearningtimeseries]. Whereas CNNs are highly effective at efficiently identifying local patterns and hierarchical structures within the input signal, the Transformer’s self-attention mechanism excels at modeling global context and capturing subtle long-range temporal dependencies.
The processing of information in our model can be viewed as a composite function operating on the input data. Let $\mathcal{X}$ represent the input, which could be potentially padded or pre-processed. The data flows through the components as follows:
  1. 1.
    CNN Feature Extractor ($\mathcal{F}$): $\mathcal{X}$ is first processed by the CNN feature extractor ($\mathcal{F}$) which maps the time series to a sequence of latent feature vectors ($L$). The feature extractor reduces the temporal dimension while increasing the feature dimensionality, capturing local patterns in the signal and potentially making the training faster due to working with shorter sequences.
  2. 2.
    Transformer Encoder (T): The latent sequence of feature vectors ($L$) is first linearly projected to match the transformer’s input dimensionality, to be followed by the addition of positional encodings to inject information about the position of each element in the sequence. To facilitate sequence-level aggregation, a special token, often denoted as [CLS], is also prepended to the sequence before the addition of positional encodings. This approach, popularized by models like BERT [bert], allows the model to learn an aggregate representation of the entire input sequence into the output state corresponding to this [CLS] token. The encoder then processes this augmented sequence ($L^{\prime}$) using stacked layers of self-attention and feed-forward networks to capture global dependencies and contextual relationships within the data, outputting an encoded sequence representation ($E$). We write,
    $E=\mathcal{T}(L^{\prime}).$ (6)
  3. 3.
    Aggregation (Agg): The aggregation step involves extracting the final hidden state vector from the transformer encoder’s output sequence $E$ that corresponds to the position of the prepended [CLS] token. This vector ($E_{agg}$) carries the aggregated information from the entire sequence and is suitable for downstream tasks such as classification or source parameter inference. Formally, we write,
    $E_{agg}=\text{Agg}(E),$ (7)
    where the function Agg specifically selects the output vector at the [CLS] token’s position.
  4. 4.
    Regression MLP Head ($\mathcal{H}$): Finally, the aggregated encoded representation $E_{agg}$ is passed through the regression MLP head ($\mathcal{H}$) which maps this fixed-size vector to e.g the predicted GW source parameters ($\mathbf{y}$) using linear layers. We write,
    $\mathbf{y}=\mathcal{H}(E_{agg}).$ (8)
In summary, the entire model can be described as a composite function where the input $\mathcal{X}$ is sequentially transformed according to
$\mathcal{X}\xrightarrow{\text{(1)}}\;\mathcal{F}(\mathcal{X})\;\xrightarrow{\text{(2)}}\;\text{concat}(\texttt{[CLS]},L)\;\xrightarrow{\text{(2)}}\;\mathcal{T}(L^{\prime})\;\xrightarrow{\text{(3)}}\;\text{Agg}(E)\;\xrightarrow{\text{(4)}}\;\mathcal{H}(E_{\text{agg}}).$ (9)
3.2 Feature extractor
The CNN feature extractor ($\mathcal{F}$), illustrated in Figure 2, is based on the InceptionTime architecture [inceptiontime], originally proposed for time-series classification. The network begins with a dedicated (stem) layer which processes the raw input signal ($\mathcal{X}$), which accepts a single-channel time series, typically of 4096 samples, using a convolution with a kernel size $k=7$ and a stride equal to 1. This convolution is followed by Batch Normalization (BN) [ioffe2015batchnormalizationacceleratingdeep] and a ReLU activation. The stem layer serves as a preparatory phase, transforming the input into an initial, richer feature representation, by projecting the input features to a higher dimensionality – we use $64$ channels as default. The core of the feature extractor is constructed from a sequence of $N_{B}$ main processing blocks stacked sequentially. Each of these main blocks internally performs a sequence of operations. First, the block processes its input through a stack of $M$ InceptionModule sub-blocks, illustrated in Figure 3. A residual connection [deepresiduallearning] is then applied around this entire stack of $M$ modules, and the result is passed through a ReLU activation. Finally, each main block is concluded by a MaxPool layer for temporal downsampling, with a kernel size and stride of 2. The latter operation reduces the sequence length (halving it at each step) and the computational cost for subsequent layers, which enables the model to learn features at broader time scales. Unlike the original InceptionTime architecture which uses Global Average Pooling for a final classification, our architecture employs this progressive MaxPool downsampling between blocks. This is an important adaptation, motivated by the computational complexity of the subsequent Transformer encoder. The self-attention mechanism in a standard Transformer has a computational and memory complexity of $O(N^{2})$, where $N$ is the sequence length. This quadratic scaling makes processing very long sequences, such as our raw GW time series, computationally prohibitive. By progressively downsampling with a stride of 2, we halve the sequence length at each step, and so reducing the computational load of the attention mechanism by a factor of four. This makes the application of a Transformer encoder more feasible while still preserving the sequential nature of the data and enabling hierarchical feature learning. While we use the standard self-attention, we note that significant research has been dedicated to mitigating this quadratic bottleneck, leading to more efficient Transformer variants with linear-time attention mechanisms [tay2022efficienttransformerssurvey].
$\mathcal{X}$ Stem $M\times$ IM 1x1C/ID +ReLU MP $N_{B}\times$ BN $L$
Figure 2: Architecture of our feature extractor ($\mathcal{F}$). The highlighted dashed box represents the main processing block, which is repeated $N_{B}$ times. IM stands for InceptionModule, and MP denotes MaxPool layer. $M$ is the number of modules stacked within each main block.
The InceptionModule follows the core idea of the Inception design, that is, parallel convolutional branches processing the input at multiple scales, with outputs concatenated along the channel dimension. Our implementation adapts the original InceptionTime formulation [inceptiontime] with a key modification to its normalization strategy. Instead of applying Batch Normalization after each internal convolutional layer, we apply a single Batch Normalization layer only after the outputs of all parallel branches have been concatenated. This design choice aims to reduce the overall computational cost of the module and provides a unified normalization step on the aggregated multi-scale feature map, which can help stabilise training by controlling the distribution of the combined features before they are passed to subsequent layers.
Following the original idea of InceptionTime block architecture [inceptiontime], we use a bottleneck layer based on a convolution with kernel $k=1$ and stride 1, which projects the channel dimension of the input time series to a smaller number. We set the default bottleneck size value to 32 output channels. This bottleneck layer significantly reduces dimensionality and model complexity, allowing for the use of larger effective filters in subsequent layers with a comparable parameter count. A ReLU activation is then applied after the bottleneck operation. The output of this path is then processed by three parallel convolutional layers with a stride of 1 and with different kernel sizes, enabling the capture of temporal patterns at different scales. For the size of the three kernels we choose $k\in\{10,20,40\}$. The use of such large kernel sizes is a key feature of the InceptionTime design, made computationally feasible by the preceding bottleneck layer. These large kernels allow the model to capture features over extended time horizons, which is particularly suitable for the potentially long-duration and complex morphologies of GW signals. We also use padding to preserve the sequence length through the module, so the output of all convolutional branches has the same sequence length. Each convolution is configured to output a common number of feature maps, $d_{branch}$, which is constant within a main block but can vary between main blocks. We moderately increase the number of filters ($d_{\text{branch}}$) compared to the design of Ref. [inceptiontime] in deeper blocks of the network. This design compensates for the reduction in temporal resolution caused by the MaxPool downsampling and allows the network to learn progressively more complex feature representations from the compressed sequence. Concurrently, a separate MaxPool pathway processes the original input ($x$), with $k=3$ and stride 1, which helps the model achieve some invariance to small temporal perturbations [inceptiontime]. The output of this MaxPooling operation is then passed through a convolution layer with kernel size $k=1$ to project its feature dimension to match the filters of the convolutional branches.
Finally, the outputs from all three parallel convolutional branches and the processed MaxPooling pathway are concatenated along their feature channel dimension and passed through a final BatchNorm (BN) layer. The entire feature extractor network leads to a final Batch Normalization layer which normalises the output feature sequence $L$. For all convolutional layers, if pretrained weights are not used, the weights are initialised using Kaiming Normal initialization [kaiming], which is well-suited for the ReLU activations used throughout the network. In our experiments, the feature extractor was configured with $N_{B}=3$ main blocks, each containing $M=3$ InceptionModules. The bottleneck size within each module was set to $d_{bottle}=32$. The number of filters per branch, $d_{branch}$, was varied for each of the main blocks, with values of $[32,48,64]$ respectively.
Input xxBottleneckReLUConvConvConvk=10k=20k=40MaxpoolConvk=1ConcatBNOutput
Figure 3: Architecture of our InceptionModule. The bottleneck layer first reduces the channel dimensionality of the input $x$. This output is then processed by multiple parallel convolutions to extract multi-scale features. Concurrently, the original input is processed by a max-pooling pathway. The outputs of all pathways are concatenated and normalised.
3.3 Transformer encoder and regression head
The resulting latent sequence of the feature extractor ($\mathcal{F}$) has dimensions $L\in\mathbb{R}^{B\times N\times d_{\mathcal{F}}}$, where $B$ is the batch size, $N$ is the sequence length reduced by $\mathcal{F}$, and $d_{\mathcal{F}}$ is the feature dimension (output channels) from $\mathcal{F}$. This sequence is first projected into the transformer’s model dimension ($d_{model}$). We use a linear projection layer with weight matrix $W_{0}\in\mathbb{R}^{d_{\mathcal{F}}\times d_{model}}$ and bias $b_{0}\in\mathbb{R}^{d_{model}}$, applied independently to each element along the sequence length dimension for all batches as
$L_{(b,i,\cdot)}=L_{(b,i,\cdot)}W_{0}+b_{0},\quad\text{for }b=1..B,i=1..N,$ (10)
which leads to the projected sequence with dimensions $L\in\mathbb{R}^{B\times N\times d_{model}}$. As described previously, a special classification token, [CLS], is prepended to $L$ forming the input sequence $L^{\prime}$, allowing the model to learn a sequence-level summary through the self-attention mechanism [bert]. Positional encodings are added to $L^{\prime}$ to provide the model with sequence order information, which the transformer itself does not inherently capture. We use standard sinusoidal encoding as in [vaswani2017attention].
Our encoder follows the standard transformer architecture [vaswani2017attention], and consists of a stack of identical layers, each comprising a multi-head self-attention mechanism and position-wise fully connected feed-forward network. Residual connections are employed around each of the two sub-layers, followed by layer normalization. GELU activation is used. The encoder processes the entire augmented sequence $L^{\prime}$ and outputs an encoded sequence $E$ of the same length, where each output vector represents its corresponding input element enriched with contextual information from the entire sequence. The output vector corresponding to the [CLS] token acts as an aggregated representation $E_{agg}$ of the whole input sequence which is then fed into the MLP regression head $\mathcal{H}$, consisting of three fully-connected linear layers, followed by GELU and dropout, except for the final output layer. $\mathcal{H}$ maps the learned, feature representation $E_{agg}$ to the inferred parameters, $y=\mathcal{H}(E_{agg})$.
Input $\mathcal{X}$ CNN Feature Extractor Transformer Encoder MLP Regression Head Output
Figure 4: Overview of our model architecture.
4 Training
4.1 Data pre-processing
For numerical stability, the original input is uniformly scaled by a factor of $10^{21}$ as, $\mathbf{x}^{\prime}_{i}=\mathbf{x}_{i}\times 10^{21}$, which transforms the typical strain amplitude of a GW closer to $\mathcal{O}(1)$. We ensure that all waveforms consist of 4096 samples, through appropriate padding. For the target parameters, $y_{i}=[m_{1},m_{2},\chi_{\rm{eff}},d]_{i}$, each component is independently normalised to [0,1] using Min-Max scaling based on the training set range. Model predictions are inverse-transformed to physical units for evaluation. We remind that distances are not sampled from an astrophysical distribution, but are instead rescaled to enforce a fixed network SNR.
The entire dataset was initially shuffled randomly to ensure an unbiased distribution of samples, and then partitioned into a training set ($D_{train}$), a validation set ($D_{val}$), and a test set ($D_{test}$) under a split ratio of 8080%, 1010% and 1010% respectively. $D_{train}$ was used for learning the model parameters $\theta$, $D_{val}$ was used to monitor the model’s performance on data not seen during training, to guide the selection of the best-performing model checkpoint, and to inform the hyperparameter optimization process. $D_{test}$ was held out during all training and model selection stages and was used at the end to provide an unbiased estimate of the final selected model’s generalization performance, including the test error and other relevant evaluation metrics.
4.2 Pretraining and training
We adopt a self-supervised pretraining strategy inspired by pretraining task proposed in [zerveas2020transformerbasedframeworkmultivariatetime]. This self-supervised objective is conceptually similar to the Masked Language Model (MLM) pretraining task popularized by BERT [bert], in that the model learns contextual representations by reconstructing a corrupted input. However, our implementation is adapted for the continuous time-series domain and differs in several key aspects. First, while BERT operates on a sequence of discrete text tokens from a finite vocabulary, our model processes a continuous-valued signal. Second, our masking strategy differs: BERT typically replaces a small percentage (usually around 15%) of individual random tokens with a special ‘[MASK]‘ token, whereas we mask contiguous spans of the time series by setting their values to zero. This span-based masking encourages the model to learn the structure of longer temporal segments. Finally, the prediction tasks are distinct: BERT performs a classification task, predicting the original discrete token ID using a Cross-Entropy loss. In contrast, our model performs a regression task, predicting a continuous latent vector ($\mathbf{z}_{\text{pred}}$) and is optimised using a MSE loss against the target latent vector ($\mathbf{z}_{\text{target}}$) derived from the uncorrupted signal. During pretraining, random contiguous segments of input waveforms are masked (set to zero), and the model is tasked with reconstructing a latent representation of the original, uncorrupted signal from this partial input. This process encourages the model to learn robust temporal dependencies and feature representations. Given an input waveform $\mathbf{x}\in\mathbb{R}^{L}$, the binary mask $\mathbf{m}\in\{0,1\}^{L_{in}}$ is generated stochastically for each sample using a process modeled as a two-state discrete-time Markov chain (see Figure 5). The two states are $S_{0}$ (being inside a masked segment, where the mask value is 0) and $S_{1}$ (being inside an unmasked segment, where the mask value is 1). The duration of stay in each state follows a geometric distribution. The probability of transitioning out of a state at any given time step is denoted by $p$. The mean segment length is therefore $1/p$. We define the mean length for masked segments as $l_{m}$ and for unmasked segments as $l_{u}$, which in turn define the transition probabilities $p_{m}=1/l_{m}$ and $p_{u}=1/l_{u}$, respectively. The process is determined by two key hyperparameters: the desired mean masked length, $l_{m}$, and the overall masking ratio, $r$, from which the mean unmasked length is derived as $l_{u}=l_{m}\cdot(1-r)/r$. The transition matrix $\mathbf{P}$ for this Markov chain is given by
$\mathbf{P}=\begin{pmatrix}1-p_{m}&p_{m}\\ p_{u}&1-p_{u}.\end{pmatrix}$ (11)
For our experiments, we set a masking ratio of $r=0.15$ and a mean masked segment length of $l_{m}=32$. This choice of a relatively long mean segment length is deliberate and tied to our network’s architecture. It serves two purposes: first, it prevents the reconstruction task from becoming trivial, as the model must learn long-range dependencies rather than simply interpolating from immediate neighbors. Second, and more critically, it ensures that the masked regions are robust to the temporal downsampling performed by the feature extractor. With a total downsampling factor of 8x in our network, a mean masked length of 32 ensures that each masked segment in the input corresponds to multiple tokens in the latent space where the reconstruction loss is computed, providing a stable and consistent learning signal. The masked vector is then applied to the input waveform via element-wise multiplication to produce the masked input
$\mathbf{x}_{m}=\mathbf{x}\odot\mathbf{m},$ (12)
and we also define a binary padding mask $\mathbf{p}\in\{0,1\}^{L}$, indicating valid input positions to distinguish from 0’s in the sequence which were due to padding. The pretraining objective involves two parallel data processing pathways. The main pathway generates a prediction from the corrupted input, while the second generates the ground truth target from the original input. The prediction path proceeds as follows: the masked waveform is passed through a feature extractor $\mathcal{F}(\cdot)$, yielding downsampled features ($\mathbf{l}_{m}\in\mathbb{R}^{B,N,d_{\mathcal{F}}}$). These are subsequently transposed and linearly projected to the transformer’s dimension $d_{model}$ by $f_{\text{proj}}(\cdot)$, on top of which positional encodings are added through $f_{\text{pos}}(\cdot)$. Schematically,
$\mathbf{l}_{m}=\mathcal{F}(\mathbf{x}_{m})\in\mathbb{R}^{C\times L^{\prime}}\;\;\to\;\;\mathbf{z}_{m}=f_{\text{proj}}(\mathbf{l}_{m}^{\top})\in\mathbb{R}^{L^{\prime}\times D}\;\;\to\;\;\mathbf{z}_{\text{pe}}=f_{\text{pos}}(\mathbf{z}_{m}).$ (13)
To prevent the self-attention mechanism from attending to padded regions, the original padding mask $\mathbf{p}$ is downsampled to the new sequence length $L^{\prime}$. This downsampling must precisely mirror the temporal reduction performed by the main feature extractor $\mathcal{F}$. We achieve this by applying a sequence of MaxPool layers (with a kernel size and stride of 2, identical to the downsamplers in $\mathcal{F}$) to the padding mask. Using Max Pooling on the binary mask correctly propagates the padding information, ensuring that a position in the resulting downsampled mask, $\mathbf{p}^{\prime}\in\{0,1\}^{L^{\prime}}$, is marked as padding (value = 0) only if all input positions in its receptive field were also padding. The resulting mask is then inverted to define the attention mask $\mathbf{m}_{\text{attn}}=\neg\mathbf{p}^{\prime}$, which is supplied to the transformer encoder. The transformer encoder $\mathcal{T}(\cdot)$ processes the input and the output is passed through a linear prediction head $f_{\text{head}}(\cdot)$ as
$\mathbf{l}_{e}=\mathcal{T}(\mathbf{z}_{\text{pe}},\mathbf{m}_{\text{attn}})\;\;\to\;\;\mathbf{z}_{\text{pred}}=f_{\text{head}}(\mathbf{l}_{e}).$ (14)
Figure 5: Illustration of the binary padding mask downsampling process. A MaxPool operation with a kernel size and stride of 2 is applied to the input mask. This ensures that a position in the downsampled mask is marked as padding (0) only if all corresponding input positions were also padding.
1 1 1 0 0 0 1 1 Input Padding Mask ($\mathbf{p}\in\{0,1\}^{L_{in}}$)1 1 0 1 Downsampled Mask ($\mathbf{p}^{\prime}\in\{0,1\}^{L_{out}}$)max(1,1)=1max(1,0)=1max(0,0)=0max(1,1)=1
The target latent representation is computed from the original (unmasked) input using the same feature extractor and projection,
$\mathbf{z}_{\text{target}}=f_{\text{proj}}(\mathcal{F}(\mathbf{x})^{\top}),$ (15)
and finally, the loss is computed only on positions that were both masked and valid (i.e., not padded). We note that the binary mask $\mathbf{m}$ is downsampled the same way as our padding mask. We define a loss mask
$\mathbf{m}_{\text{loss}}=\text{downsample}(\neg\mathbf{m}\land\mathbf{p})\in\{0,1\}^{L^{\prime}}.$ (16)
This ensures that only valid and originally masked tokens contribute to the reconstruction loss. Denoting with $I=\{(b,i)\mid\mathbf{m}_{\text{loss},(b,i)}=1\}$ the set of batch and sequence indices where the loss should be computed, pretraining aims to optimise the mean squared error (MSE) between the predicted and target latent representations,
$\mathcal{L}_{pretrain}=\frac{1}{|I|}\sum_{(b,i)\in I}\left\lVert\mathbf{z}_{\text{target},(b,i,\cdot)}-\mathbf{z}_{\text{pred},(b,i,\cdot)}\right\rVert_{2}^{2}.$ (17)
Following pretraining, the model is trained/fine-tuned to predict the target physical parameters $\mathbf{y}_{i}$ (e.g masses, spins) using supervised training. We minimize the MSE over batches $\mathcal{B}\subset D_{\text{train}}$,
$\mathcal{L}_{\mathrm{MSE}}(\theta;\mathcal{B})=\frac{1}{B}\sum_{(\mathbf{x}_{i},\mathbf{y}_{i})\in\mathcal{B}}\|\mathbf{y}_{i}-f(\mathbf{x}_{i};\theta)\|_{2}^{2},$ (18)
where $f(\mathbf{x}_{i};\theta)$ is the predicted target vector and $\theta$ the deep model’s inherent parameters. Fine-tuning uses the AdamW optimiser [loshchilov2019decoupledweightdecayregularization] with learning rate $5e^{-4}$, cosine scheduler with linear 10% warmup, batch size $B=250$, and a maximum of $E=120$ epochs. Gradient clipping with threshold 1.0 is applied to prevent instability. The hyperparameters of our model are summarized in Table 10. We evaluate the trained model, which we denote as $f(\mathbf{x};\theta^{**})$, on the test dataset ($D_{\text{test}}$) using two regression metrics for each predicted physical parameter. Let us denote with $y_{ij}$ and $\hat{y}_{ij}$ the true and predicted values of the $j$-th parameter for the $i$-th test sample after inverse scaling, and with $\bar{y}_{j}$ the mean of the true $j$-th parameter over $D_{\text{test}}$ respectively. We then define the metrics
$\mathcal{L}_{\mathrm{MAE}_{j}}=\frac{1}{n}\sum_{i=1}^{n}|y_{ij}-\hat{y}_{ij}|,\;\;\mathcal{L}_{R^{2}_{j}}=1-\frac{\sum_{i=1}^{n}(y_{ij}-\hat{y}_{ij})^{2}}{\sum_{i=1}^{n}(y_{ij}-\bar{y}_{j})^{2}}.$ (19)
The Mean Absolute Error (MAE) captures average absolute errors, while $R^{2}_{j}$ measures the proportion of variance explained by the model. Perfect predictions yield $R^{2}_{j}=1$, while $R^{2}_{j}<0$ indicates worse performance than predicting the mean.
4.3 Model uncertainty
For uncertainty estimation we employ Monte Carlo (MC) Dropout [gal2016dropoutbayesianapproximationrepresenting] as a computationally efficient method to perform probabilistic inference and estimate the predictive uncertainty of our deep learning model. In this context, it is useful to distinguish between two primary types of uncertainty [kendall2017uncertaintiesneedbayesiandeep]: epistemic uncertainty, which captures our ignorance about the model parameters and can be reduced with more data, and aleatoric uncertainty, which captures the inherent noise or randomness in the data itself. The key idea behind MC Dropout is to retain dropout during inference, not just training. Given an input $\mathbf{x}_{i}$, we perform $T$ stochastic forward passes through the trained network $f(\cdot;\theta)$, applying different dropout masks in each pass. This yields a set of $T$ output samples,
$\left\{\hat{\mathbf{y}}_{i}^{(t)}=f(\mathbf{x}_{i};\theta^{(t)})\right\}_{t=1}^{T}.$ (20)
This represents a set of $T$ stochastic predictions for input $\mathbf{x}_{i}$, where $\theta^{(t)}$ denotes the parameters randomly thinned by the $t$-th dropout mask. These samples approximate the posterior predictive distribution over $\mathbf{y}_{i}$, allowing us to estimate both the expected prediction and the associated uncertainty. The point estimate for the target vector $\hat{\mathbf{y}}_{i,MC}$ is taken as the sample mean of the ensemble, and the uncertainty for each individual output parameter component $y_{ij}$ (the $j$-th component of $\mathbf{y}_{i}$) is estimated using the sample variance of the $T$ predictions for that specific component as
$\hat{\mathbf{y}}_{i,\text{MC}}=\frac{1}{T}\sum_{t=1}^{T}\hat{\mathbf{y}}_{i,t},\;\;\;\widehat{\text{Var}}(\hat{y}_{ij})=\frac{1}{T-1}\sum_{t=1}^{T}(\hat{y}_{ij,t}-\hat{y}_{ij,\text{MC}})^{2}.$ (21)
The square root of this variance, $\hat{\sigma}(\hat{y}_{ij})=\sqrt{\widehat{\text{Var}}(\hat{y}_{ij})}$ is the standard deviation for each parameter component, which serves as our primary measure of model uncertainty. The sample variance calculated via MC Dropout serves as an approximation of the total predictive uncertainty. It implicitly combines both the epistemic uncertainty, arising from the stochasticity of the dropout masks, which approximates uncertainty over the model’s weights, and the aleatoric uncertainty that is propagated through the deterministic parts of the network from the input data [kendall2017uncertaintiesneedbayesiandeep]. For a computation of the Bayesian posterior which captures all kinds of uncertainty, an approach such as simulation based inference would be appropriate.
5 Results
Here, we present our results for the parameter inference of noisy waveforms and overall performance of our model. To understand the computational requirements and training dynamics, we trained our model on a small ($\sim 20,000$), intermediate ($\sim 50,000$), and large ($\sim 80,000$) dataset. The train/validation/testing sets were created using our strategy described in the Section 4.1, whereas the small and intermediate ones taken as subsets of the large training set. For all experiments the validation and test sets remained fixed to ensure that the performance improvements observed are attributable to the increase in training data, not variations in the evaluation sets. We notice that all training and evaluation were performed on a single NVIDIA A100 GPU.
5.1 Effect of pretraining
The results associated to the effect of pretraining at fixed detector noise (ET) are summarised in Tables  1 - 3. Tables 1 and 2 summarise the effect of pretraining on the convergence speed and performance metric respectively, comparing training from scratch and training/fine-tuning on top of a pretrained model. The relative performance gains are shown in Table 3. For each combination of dataset size and signal-to-noise ratio (SNR), the “Target Loss” corresponds to the best validation loss achieved by a model trained from scratch (no pretraining). The results clearly show that pretraining substantially accelerates convergence across all SNRs and dataset sizes, with speedup factors ranging from approximately $1.4\times$ to an impressive speedup of $6.6\times$. The most pronounced benefits of pretraining are clearly observed for small and intermediate datasets, where convergence can be more than $6\times$ faster, such as for SNR = 10 in the small and intermediate dataset scenarios. Remarkably, the speedup advantage is generally greater at lower SNRs, indicating that pretraining is especially beneficial when the signal is weaker and the learning task is more challenging. For large datasets, the speedup effect is less pronounced, particularly at SNR = 10, though the reduction in training time remains sizable. At higher SNR with a large dataset, the speedup is still notable, suggesting that while having a large amount of data partially compensates for the absence of pretraining, the use of pretrained models continues to offer substantial gains. Overall, pretraining in GraviBERT is highly effective for reducing training times and enables more rapid model development, which is particularly valuable in computationally intensive experiments.
Pretraining enhances also the performance metrics of the model. The absolute performance metrics for the pretrained and fine-tuned models are summarized in Table 2, while the fractional gains compared to training from scratch in Table 3. The metrics are computed as the average over 100 stochastic forward passes per sample using MC Dropouts based on the test set. Across all evaluated cases, the pretrained/fine-tuned model consistently outperforms the model trained from scratch, achieving lower MAE for all parameters and, in some cases, higher $R^{2}$ values although the latter is typically mildly affected. At fixed dataset size, pretraining significantly improves MAE with increasing SNR, with the biggest effect in the large dataset where the percentage MAE improvement is about $3\times$ larger. For the most challenging case, specifically the case of SNR = 10 and the small dataset, the benefits of pretraining are pronounced, with a MAE reduction ranging between $6\%-14\%$ across parameters. The best reduction is found for the case of the large dataset and SNR = 30, ranging between $11\%-30\%$. This can be attributed to the sufficiently large amount of data in the large dataset combined with the high SNR which allow the model to adopt and train better. One notices that the intermediate dataset shows suboptimal gains compared to both smaller and larger datasets. This appears to be due to the dataset being large enough for the model to start overfitting to the training data, but not large enough to fully benefit from the pretrained representations or to learn robust patterns from scratch. Another possibility is that the pretrained features may compete with task-specific learning in a suboptimal way leading to this behavior. Furthermore, this could also be (partially) attributed to the use of a single set of hyperparameters across all experiments. The optimal hyperparameters for fine-tuning a pretrained model, which often benefit from a smaller learning rate, may differ from those for training a model from scratch. A more exhaustive study would involve performing a separate hyperparameter search for each condition. However, our results using a single, tuned configuration consistently demonstrate a network positive benefit from pretraining across all dataset sizes. We notice that for all the above cases $R^{2}\geq 0.90$ as shown in Table 2.
Understanding how model performance scales with dataset size informs future data requirements and expected performance gains  [kaplan2020scalinglawsneurallanguage]. The test loss (measured on models trained from scratch) follows a power-law trend as a function of training dataset size $D$,
$L(D)\propto D^{-\alpha_{D}},\quad\text{with}\quad\alpha_{D}=\begin{cases}0.3701,&\text{for SNR = 10}\\ 0.3624,&\text{for SNR = 30},\end{cases}$ (22)
suggesting GraviBERT’s performance continues to improve with more data, while the scaling exponent varies extremely mildly across both the SNRs studied.
5.2 Performance on parameter inference
[…]
Figure 8 shows the precision of the model’s predictions with respect to the target values across our grid. For the component masses, we evaluate the mean relative precision across our predictions, defined as $100\times(\text{prediction}-\text{target})/\text{target}$. For the effective spin, we report the mean of the absolute difference $(\text{prediction}-\text{target})$, in order to avoid amplification due to division by values near zero. The mean relative precision across the sample for $m_{1}$ is $6\%$ at SNR = 10 and $3.3\%$ at SNR = 30. The similar values for $m_{2}$ are $3.4\%$ and $1.1\%$, while for distance $8.3\%$ and $4.1\%$ respectively. For $\chi_{\rm{eff}}$, the absolute mean error value is $3.8\times 10^{-3}$ at SNR = 10 and $6\times 10^{-5}$ at SNR = 30. To enable an approximate comparison with real observations, we focus on model predictions at SNR = 10, and select the event GW190828_065509 from the third Gravitational Wave Transient Catalog (GWTC-3) [Abbott2023gwtc3] of the LIGO-Virgo (LV) Collaboration, which has an SNR compatible with $10$  [^3] The reported central values for masses and effective spin are $m_{1}=23.7M_{\odot},m_{2}=10.4M_{\odot},\chi_{\rm{eff}}=0.05$.. To compare our model’s mean relative errors with the $90\%$ credible intervals reported by the LV Collaboration, we convert them to equivalent $90\%$ errors under the assumption of Gaussianity, which amounts to multiplying them by $\simeq 1.645$. For $\{m_{1},m_{2},\chi_{\rm{eff}},d\}$ we find for the average $90\%$ error across our sample $\{9.9\%,5.7\%,0.006,13.6\%\}$, while the corresponding errors reported by the LV Collaboration for the event GW190828_065509 are $\{28.5\%,28.9\%,0.165,44\%\}$. Notice that we have symmetrised the observation errors to facilitate comparison, i.e $\sigma_{\rm 90\%}=(\sigma_{\rm 90\%upper}+\sigma_{\rm 90\%lower})/2$. These numbers should not be interpreted as a direct posterior-to-posterior comparison, and serves only as a back-of-the-envelope estimate. A Bayesian analysis would be needed for a proper comparison. We also emphasise that our distance refers to an effective luminosity distance rescaled to fixed SNR, not an actual astrophysical one.
5.3 Domain adaptation
Domain adaptation refers to transferring knowledge across domains and is central to foundation-style models, which learn general rather than domain-specific representations. Effective adaptation indicates that the model has captured fundamental structures in the data, with the pretraining process providing robust, transferable features which subsequently require only light fine-tuning. In our case, this invariance is assessed only within the shared ET–LV frequency band ($\geq 20$ Hz). We focus on transferring from ET to LV noise, evaluating whether the model can adopt detector-agnostic GW features rather than memorising detector-specific noise.
We first evaluate zero-shot inference by applying the model trained purely on ET noise directly to the LV one, without any fine-tuning or weight updates. This results in a loss of predictive power, indicating that ET-trained features do not transfer to the LV domain in a zero-shot setting. This is naturally explained by the model having learned a detector-conditioned latent space rather than an invariant one, and therefore, requiring domain adaptation through fine-tuning. Implementing transfer learning through fine tuning, the performance gain is substantial compared to training from scratch across all inferred parameters, ranging between $9\%-47\%$ improvement for MAE (see Tables 4 and 5). The spin parameter shows the biggest improvements ($21-47\%$), suggesting that the pretrained model learned particularly robust representations of spin-related features that transfer effectively between detectors. The mass parameters and distance also show consistent improvements of $13-27\%$, indicating that the model successfully captures detector-agnostic physical relationships. The convergence speed results show that the pretrained model achieves the same performance targets as models trained from scratch in dramatically less time, up to $15\times$ faster. This speedup is particularly pronounced at lower SNR = 10, where the pretrained model reaches target performance in just 2-3 minutes compared to 30-50 minutes for scratch training. Interestingly, the speedup is somewhat reduced SNR = 30, possibly because cleaner signals allow scratch-trained models to converge more efficiently. These results collectively demonstrate that pretraining on one detector configuration creates transferable representations that not only improve final performance, but also significantly reduce the computational cost of adapting to new detector environments. We remind that while fine-tuning on LV noise shows successful adaptation, both detectors still share broadly similar noise structure above 20 Hz. The more demanding task of adapting LV-trained models to ET’s full 5 Hz range remains open for future work.
To evaluate the generalisation of GraviBERT against waveform modelling, we performed a cross-approximant validation. We first reconstructed the held-out test samples from the large dataset (originally generated with SEOBNRv4) using the IMRPhenomD approximant, adjusting the signals to match the temporal grid of the original SEOBNRv4 training data. Inference was then performed on these new waveforms using the frozen weights of our fully trained model (pretrained and trained on ET noise and SEOBNRv4 approximant) without any further fine-tuning, that is, at zero-shot inference. The results are shown in Table 6. While the baseline model achieves $R^{2}>0.93$ for all parameters at SNR = 10, the cross-approximant inference appears to introduce a systematic bias. This effect is particularly evident for the effective spin parameter, where $R^{2}=0.65$ at SNR = 10. Spin is intimately connected with the subtle phase evolution during inspiral, which is modelled differently in SEOBNRv4 and IMRPhenomD. Nevertheless, the model retains predictive power for the dominant parameters $(m_{1},m_{2},d)$ at zero-shot inference, with $R^{2}$ remaining consistently between [0.80,0.97]. This suggests that the model has extracted representations of the signal envelope and frequency chirp, which are approximant-agnostic, while the specific phase alignment remains sensitive to the training domain.
[…]
5.4 Benchmarking
To contextualise our results, we compare GraviBERT with other approaches for gravitational wave parameter estimation. Since we focus on point estimates in this work, we compare with methods using similar inference targets rather than full neural posterior estimation frameworks. We should emphasise that direct comparisons with prior ML methods are complicated by differences in parameter dimensionality, waveform models, noise assumptions, signal duration, architectures, and hardware platforms. Several works focus on predicting only two intrinsic parameters (e.g., chirp mass and mass ratio), while others use neural density estimation to approximate full posterior distributions. Our model performs four-parameter regression $(m_{1},m_{2},\chi_{\mathrm{eff}},d)$ with a hybrid CNN-Transformer architecture designed for representation learning and transferability. Most importantly, we have not performed a hyperparameter optimisation which could significantly improve the performance. Numerical comparisons should therefore be interpreted as contextual indicators rather than direct benchmarks.
George & Huerta (2018) [george2018deep2_inference] demonstrated CNN-based parameter estimation achieving mean relative errors below 10% for mass parameters on real LV data. Shen et al. (2022) [shen2022deep] achieved 2 ms inference time using Bayesian neural networks trained on over 10 million waveforms for 4-parameter estimation, with relative errors of 7–12% for masses and 5–13% for remnant properties at SNR $\geq$ 15.
[…]
We remind that we did not perform hyperparameter optimisation, as our aim was to introduce the architecture and demonstrate that (i) BERT-style masked pretraining greatly speeds up convergence and improves accuracy, (ii) pretrained models enable strong transfer learning, and (iii) hybrid CNN-Transformer models capture the essential features of the GW signals studied in this work. These strategies support both point estimation and Bayesian inference.
6 Conclusions and Outlook
We have introduced GraviBERT, a novel and versatile model for parameter inference of GW time-series, through the fusion of an Inception-inspired multi-scale convolutional frontend with a transformer encoder. GraviBERT infers source parameters, such as masses and spins, directly from noisy time-domain waveforms down to few-percent level, without the need for explicit denoising steps. Our two-stage training strategy, leveraging BERT-style pretraining followed by targeted fine-tuning, yields substantial improvements in both training efficiency and inference accuracy, and enables adaptation to new detector configurations and waveform approximants. When evaluated on in-domain data (ET), pretraining reduces the mean absolute error by up to $30\%$ and accelerates training convergence by up to $6\times$. Furthermore, the mean relative precision of the inferred masses and distances reaches the few-percent level, while the mean absolute error in the effective spin is as low as $\sim 10^{-3}$ for the challenging case of SNR = 10  [^4] We notice that these results were derived without a thorough hyperparameter study for our model.. We remind that our distance refers to an effective luminosity distance rescaled to fixed SNR, not an actual astrophysical one.
The advantages are even more pronounced in domain adaptation from ET to LV data. By fine-tuning on a small target dataset on LV data, the pretrained model on ET noise converges up to $15\times$ faster and achieves an error reduction of up to $40\%$ compared to a model trained from scratch on LV. This enhanced predictive precision is particularly notable in the low SNR regime, while the flexibility of the model’s architecture paves the way for broad applicability across multi-messenger astronomy. We note that while our estimation of inherent model uncertainty relied on the probabilistic approach of MC Dropouts, our parameter inference itself was not Bayesian. A Bayesian treatment could be pursued by replacing the regression head with normalising flows.
[…]
The significance of developing foundation-style models lies in their ability to generalise across a wide range of tasks and datasets, serving as adaptable machines for diverse downstream applications. Foundation models have transformed fields such as natural language processing enabling rapid fine-tuning and transfer learning. For GW astrophysics, they promise a similar leap: accelerate scientific discovery by enabling fast, accurate, and robust inference as new detectors, noise environments, and physical scenarios emerge. In this regard, GraviBERT represents an important step forward towards establishing a flexible, pretrained baseline framework for multi-messenger astronomy, which can be subsequently fine-tuned efficiently.
Exciting avenues for future work include applying GraviBERT to test subtle deviations from GR in gravitational waveforms, as well as probing new physics—such as dark energy—through its impact on the waveform’s amplitude and phase evolution [Berti2015, Saltas:2018fwy, Abac:2025saz, LISA:2022kgy, Barausse:2020rsu]. A critical next step is to extend our analysis to fully Bayesian parameter estimation, and to enable inference of the complete set of source parameters for binary mergers. Additionally, our modelling can be expanded to cover neutron star mergers, extreme mass ratio inspirals, alternative waveform approximants, and predicting the time to merger from inspiral waveforms. A promising direction is also the extension of GraviBERT to multi-messenger inference by incorporating electromagnetic and neutrino signals. On the architectural front, enhancements such as frequency-domain inputs and comprehensive hyperparameter optimisation could significantly improve performance. Together, these developments will help establish GraviBERT as a robust multi-modal framework for multi-messenger astrophysics.
Acknowledgements
We thank François Charton, Nikos Karnesis, Fillipos Kokkinos, Nick Stergioulas and David Trestini for discussions and feedback. This work was funded by the Czech Grant Agency (GAČR) under the grant number 21-16583M. The bulk of our computations were carried out at the MetaCentrum cluster in Czechia.
Table 1: Comparison of convergence speed between models trained from scratch and pretrained models at fixed detector configuration (ET). For each condition, the “Target Loss" is the best validation loss achieved by the corresponding model trained from scratch. The speedup factor quantifies how faster the pretrained model reaches this same performance target. Results are discussed in Section 5.
SNR Target Loss Time (Scratch) Time (Pretrained) Speedup
(Best Val Loss) (minutes) (minutes)
Dataset Size: Small
SNR 10 0.005474 79.64 13.05 6.10
SNR 30 0.001822 98.61 20.86 4.73
Dataset Size: Intermediate
SNR 10 0.004029 196.07 29.75 6.59
SNR 30 0.001299 222.02 51.51 4.31
Dataset Size: Large
SNR 10 0.003172 242.82 169.63 1.43
SNR 30 0.000952 348.69 87.92 3.97
    Time is measured as the cumulative training time until the validation loss first drops below the corresponding target loss for that specific SNR and dataset size.
Table 2: Per-parameter metrics across different dataset sizes and SNR values for the model trained from scratch (not pretrained) and the pretrained and fine-tuned model, evaluated on the held-out test set with size of $10,000$ samples. The same detector configuration is kept throughout (ET). The reported metrics are measured on the mean values obtained from an ensemble of $100$ stochastic forward passes using MC Dropout for each test sample. We remind that $d$ corresponds to an effective luminosity distance rescaled to fixed SNR.
(a) Trained from Scratch
MAE
Parameter SNR 10 SNR 30 SNR 10 SNR 30
Dataset Size: Small
$m_{1}$ 8.51 4.79 0.90 0.97
$m_{2}$ 6.97 4.24 0.93 0.97
$\chi_{\rm{eff}}$ 0.05 0.03 0.86 0.96
$d$ 7401.28 1238.39 0.96 0.99
Intermediate
$m_{1}$ 7.21 4.19 0.92 0.97
$m_{2}$ 6.18 3.80 0.94 0.98
$\chi_{\rm{eff}}$ 0.04 0.02 0.91 0.98
$d$ 6232.91 1025.78 0.96 0.99
Large
$m_{1}$ 6.02 3.49 0.94 0.98
$m_{2}$ 5.10 3.24 0.95 0.98
$\chi_{\rm{eff}}$ 0.03 0.02 0.93 0.98
$d$ 5414.76 867.37 0.96 0.99
(b) Pretrained + Fine-tuned
MAE
Parameter SNR 10 SNR 30 SNR 10 SNR 30
Dataset Size: Small
$m_{1}$ 7.80 4.25 0.91 0.97
$m_{2}$ 6.58 3.82 0.93 0.98
$\chi_{\rm{eff}}$ 0.04 0.02 0.90 0.97
$d$ 6807.13 1144.08 0.95 0.99
Intermediate
$m_{1}$ 7.06 4.01 0.92 0.97
$m_{2}$ 5.98 3.69 0.94 0.98
$\chi_{\rm{eff}}$ 0.04 0.02 0.92 0.98
$d$ 6154.66 986.82 0.96 0.99
Large
$m_{1}$ 5.44 2.44 0.94 0.99
$m_{2}$ 4.51 2.23 0.96 0.99
$\chi_{\rm{eff}}$ 0.03 0.02 0.93 0.98
$d$ 5043.25 639.72 0.96 0.99
Table 3: Relative MAE improvement from pretraining followed by fine-tuning, compared to training from scratch, according to the results of Table 2. Positive values indicate the percentage reduction in error achieved through pretraining.
Performance Gain on MAE (%)
Parameter SNR 10 SNR 30
Dataset Size: Small
$m_{1}$ $\approx 8\%$ $\approx 11\%$
$m_{2}$ $\approx 6\%$ $\approx 10\%$
$\chi_{\rm{eff}}$ $\approx 14\%$ $\approx 5\%$
$d$ $\approx 8\%$ $\approx 8\%$
Dataset Size: Intermediate
$m_{1}$ $\approx 2\%$ $\approx 4\%$
$m_{2}$ $\approx 3\%$ $\approx 3\%$
$\chi_{\rm{eff}}$ $\approx 4\%$ $\approx 7\%$
$d$ $\approx 1\%$ $\approx 4\%$
Dataset Size: Large
$m_{1}$ $\approx 10\%$ $\approx 30\%$
$m_{2}$ $\approx 11\%$ $\approx 31\%$
$\chi_{\rm{eff}}$ $\approx 3\%$ $\approx 11\%$
$d$ $\approx 7\%$ $\approx 26\%$
Table 4: Domain adaptation performance of the pretrained model evaluated with MAE. The comparison is between the model pretrained on ET noise and fine-tuned on LV noise waveforms, and a model trained on LV noise waveforms from scratch. The clean waveforms used cross both datasets were generated using the same procedure, with only the detector noise differing between domains.
SNR 10 SNR 30
Fine-tuning Parameter MAE MAE Perf. MAE MAE Perf.
Dataset Size (Scratch) (Fine-tuned) Gain (Scratch) (Fine-tuned) Gain
5,000 Samples $m_{1}$ 12.54 9.22 $\approx 26\%$ 6.39 4.91 $\approx 23\%$
$m_{2}$ 9.42 7.75 $\approx 18\%$ 5.64 4.41 $\approx 22\%$
$\chi_{\rm{eff}}$ 0.095 0.05 $\approx 43\%$ 0.05 0.03 $\approx 47\%$
$d$ 543.63 406.15 $\approx 25\%$ 87.85 63.82 $\approx 27\%$
10,000 Samples $m_{1}$ 10.43 8.64 $\approx 17\%$ 5.80 4.79 $\approx 18\%$
$m_{2}$ 8.18 7.12 $\approx 13\%$ 5.02 4.36 $\approx 13\%$
$\chi_{\rm{eff}}$ 0.07 0.05 $\approx 21\%$ 0.04 0.03 $\approx 32\%$
$d$ 448.28 406.89 $\approx 9\%$ 72.77 58.81 $\approx 19\%$
Table 5: Same as Table 4, but comparing the convergence speed between the model pretrained on ET noise and fine-tuned on LV noise waveforms, and a model trained on LV noise waveforms from scratch. As before, the “Target Loss" is the best validation loss achieved by the corresponding model trained from scratch and the speedup factor quantifies how many times faster the pretrained model reaches this same performance target.
SNR Target Loss Time (Scratch) Time (Pretrained) Speedup
(Best Val Loss) (minutes) (minutes)
Dataset Size: 5,000 Samples
SNR 10 0.012263 31.97 2.12 15.05
SNR 30 0.004212 29.65 5.13 5.78
Dataset Size: 10,000 Samples
SNR 10 0.008057 52.17 3.42 15.24
SNR 30 0.002486 46.97 7.79 6.03
    Time is measured as the cumulative training time until the validation loss first drops below the corresponding target loss for that specific SNR and dataset size.
Table 6: Zero-shot performance of the model on a new waveform approximant at fixed detector noise (ET), quantified using MAE and $R^{2}$. The model was pretrained and subsequently trained on SEOBNRv4. It can be seen that zero-shot inference (without fine-tuning) already yields encouraging performance in the cross-approximant setting (see also Section 5.3).
MAE
Parameter SNR 10 SNR 30 SNR 10 SNR 30
$m_{1}$ 11.16 6.44 0.80 0.93
$m_{2}$ 9.27 6.67 0.86 0.92
$\chi_{\rm{eff}}$ 0.08 0.07 0.65 0.70
$d$ 10320.96 1840.13 0.87 0.97
Table 7: Domain adaptation performance of the pretrained model on a new waveform approximant at fixed detector noise (ET), evaluated with MAE. The comparison is between the model pretrained on SEOBNRv4 waveforms and fine-tuned on IMRPhenomD waveforms, and a model trained on IMRPhenomD waveforms from scratch.
SNR 10 SNR 30
Fine-tuning Parameter MAE MAE Perf. MAE MAE Perf.
Dataset Size (Scratch) (Fine-tuned) Gain (Scratch) (Fine-tuned) Gain
5,000 Samples $m_{1}$ 13.85 8.80 $\approx 36\%$ 6.36 4.64 $\approx 27\%$
$m_{2}$ 9.51 6.91 $\approx 27\%$ 5.42 4.10 $\approx 24\%$
$\chi_{\rm{eff}}$ 0.091 0.055 $\approx 39\%$ 0.06 0.03 $\approx 44\%$
$d$ 11420.42 7471.91 $\approx 35\%$ 1839.72 1291.50 $\approx 30\%$
10,000 Samples $m_{1}$ 11.44 8.19 $\approx 28\%$ 5.46 4.29 $\approx 21\%$
$m_{2}$ 8.14 6.48 $\approx 20\%$ 4.56 3.70 $\approx 19\%$
$\chi_{\rm{eff}}$ 0.068 0.049 $\approx 28\%$ 0.039 0.028 $\approx 29\%$
$d$ 9371.46 6793.89 $\approx 27\%$ 1454.00 1095.89 $\approx 25\%$
Table 8: Same as Table 7, but comparing the speedup gain. The “Target Loss" is the best validation loss achieved by the corresponding model trained from scratch and the speedup factor quantifies how faster the pretrained model reaches the same performance target.
SNR Target Loss Time (Scratch) Time (Pretrained) Speedup
(Best Val Loss) (minutes) (minutes)
Dataset Size: 5,000 Samples
SNR 10 0.014031 34.12 2.24 15.21
SNR 30 0.004849 30.37 3.52 8.63
Dataset Size: 10,000 Samples
SNR 10 0.009436 52.46 3.65 14.36
SNR 30 0.002799 53.36 8.29 6.44
    Time is measured as the cumulative training time until the validation loss first drops below the corresponding target loss for that specific SNR and dataset size.
Table 9: Inference latency comparison for different batch sizes for the current version of GraviBERT, without hyperparameter optimisation. We consider a standard forward pass (no MC Dropout activated), and one with MC Dropout turned on. The results are discussed in Section 5.4.
Batch Size Standard (ms) MC Dropout (ms) Overhead (%) Throughput (samples/s)
1 $10.6\pm 0.5$ $10.4\pm 0.5$ -2.2 96.3
8 $11.0\pm 0.5$ $10.8\pm 0.3$ -2.0 743.9
32 $18.8\pm 0.2$ $20.1\pm 0.2$ 7.1 1590.2
64 $33.8\pm 0.3$ $36.8\pm 0.2$ 8.8 1741.1
128 $62.9\pm 0.3$ $69.5\pm 0.2$ 10.4 1843.0
250 $119.4\pm 0.3$ $132.0\pm 0.3$ 10.5 1894.5
[…]
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption SNR 10SNR 30Parameter $m_{1}$Parameter $m_{2}$Parameter $\chi_{\rm eff}$Parameter $d$
Figure 7: Same as Figure 6, but for the model which was pretrained and subsequently fine-tuned on the large dataset.
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption SNR 10SNR 30Parameter $m_{1}$Parameter $m_{2}$Parameter $\chi_{\rm{eff}}$Parameter $d$
Figure 8: Prediction errors across our grid sample, for the component masses, the effective spin, the distance and for both fixed-SNR cases of our work. For all parameters except the effective spin, the errors are shown as relative differences with respect to the corresponding target values, (prediction - target)/target. For the effective spin, only the actual difference is used to avoid numerical instability from dividing by small values. The dashed lines represent the mean of the respective values across the sample. We discuss an approximate comparison of these results with the typical errors of real waveform inferences at the end of Section 5.
Table 10: Hyperparameter configuration for the pretraining and training stages of the model.
Stage / Component Parameter Value
Pretraining
Optimiser AdamW
Learning Rate $1\times 10^{-4}$
LR Scheduler Cosine Annealing with w/ Warmup
Warmup Ratio 0.1
Batch Size 250
Max Epochs 200 (with saving the best Val checkpoint)
Masking Ratio ($r$) 0.15
Mean Masked Length ($l_{m}$) 32
Fine-tuning and Training from scratch
Training Optimiser AdamW
Learning Rate $5\times 10^{-4}$
Weight Decay $1\times 10^{-2}$
LR Scheduler Cosine Annealing w/ Warmup
Warmup Ratio 0.1
Batch Size 250
Max Epochs 120 (with saving the best Val checkpoint)
Feature Extractor Initial Stem Channels 64
Num. Main Blocks ($N_{B}$) 3
Modules per Block ($M$) 3
Bottleneck Channels 32
Kernel Sizes [10, 20, 40]
Filters per Branch ($d_{branch}$) [32, 48, 64]
Transformer Encoder Model Dimension ($d_{model}$) 256
Num. Layers 4
Num. Heads ($n_{heads}$) 4
Feed-forward Dim. ($d_{ff}$) 512
Dropout Rate 0.10
MLP Head Num. Hidden Layers 2
Hidden Dims [256, 64]
Output layer Dim 4
Dropout Rate 0.10